OpenStudy (anonymous):

The body mass index,I, can be used to determine an individual's risk for heart disease. An index less than 25 indicate a low risk. The body mass idex is given by the formula, or model, I=700w/H^2, where w=weight, in pounds, and h= height in inches. Francis weights 163 pounds and stands 69 iches tall. What is his approximate body mass index? Find and inequality describing all weights W that Franis can have and be in the low risk category.

OpenStudy (gw2011):

Body Mass Index: I=[(700)(163)]/(69)^2 I=114,100/4761 I=23.966 I=24 approximately Inequality: 700w/H^2<25 700w<(25)(H^2) w<25H^2/700 w<H^2/28